OpenStudy (anonymous):

Find the coordinates for the midpoint of the segment with endpoints given. P1 P2 M (-16, 0) (0, -16) ( , )

OpenStudy (anonymous):

(-8,-8)

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

(x1+x2/2), (y1+y2/2)......x = (-16+0 /2) = -8 ...y=(0+-16/2)=-8 ... so the midpoint is (-8,-8)